# Optomechanical self-organization in a mesoscopic atom array

### v1

@@ Abstract
A fundamental challenge in modern physics is bridging the gap between an exact description of few-body microscopic systems and the emergent description of many-body macroscopic systems. This gap may be bridged by tracing the properties of well-controlled intermediate-size mesoscopic systems. Here, we study mesoscopic signatures of a spatial self-organization phase transition in deterministically prepared arrays of between 10 and 22 atoms inside an optical cavity. Through precise engineering of the atom-cavity interactions, we reveal how critical behavior depends on atom number. We identify characteristic dynamical behaviors related to symmetry breaking and system size in the self-organized regime, and observe a finite optomechanical susceptibility at the critical point. This work opens the door to probing particle-number- and time-resolved properties of phase transitions in mesoscopic systems.
@@
The scale of a physical system plays a key role in determining its material properties. With increasing size comes increasing complexity, and complexity leads to the emergence of macroscopic phenomena that, often, cannot be extrapolated from microscopic behaviors [1]. A well-known example of this paradigm is the appearance of phase transitions in the thermodynamic limit, i.e. as a system’s particle number tends to infinity. In between the microscopic and macroscopic extremes lies the mesoscopic regime. Mesoscopic systems not only exhibit characteristics of macroscopic systems, but also host a variety of rich properties themselves. In systems of intermediate size, assumptions of local equilibrium break down and internal microstructures may affect properties on longer length scales [2]. The fact that microscopic fluctuations have a measurable influence on mesoscopic scales also makes mesoscopic systems an interesting playground for the study of critical phenomena and phase transitions. Predicted behaviors of mesoscopic systems near critical points include the smoothing out of the slope of the order parameter [3], finite-size scaling exponents [4], and switching dynamics between symmetry-broken phases [5, 6, 7].
@@
Technical advances in engineering, probing and measuring mesoscopic systems have propelled experiments in a variety of physical platforms. Notable examples include solid state systems [8, 9], nanostructured [10] and nanomechanical systems [11, 12, 13, 14], nanomagnets [15, 16], and cold atoms and ions [17, 18, 19, 20, 21, 22, 23]. Breakthroughs in the control of atoms trapped in optical tweezer arrays represent a leap forward in precisely constructing mesoscopic, ultracold atomic systems. We show that deterministic control over particle number and interaction strength can be achieved with an optical tweezer array inside a standing-wave optical cavity. The unprecendented versatility in the positioning of individual atoms offered by tweezer arrays has enabled numerous advances in quantum simulation, sensing, and information processing [24, 25]. We use the tweezers to adjust the position of each atom with respect to the spatial structure of the cavity field, thereby precisely programming the atom-cavity coupling strength. This system uniquely lends itself to the detailed study of optomechanics in mesoscopic samples.
@@ I Cavity-Mediated interactions and Self-organization
[…]
@@
To realize mesoscopic self-organization, we use the experimental setup schematically shown in Fig. 1A. It contains a one-dimensional tweezer array into which we load a deterministic number $N$ of 87Rb atoms. The array is aligned along the axis of a high-finesse Fabry-Pérot optical cavity (the $z$-axis). The cavity is placed in-vacuum and has a TEM00 mode whose frequency, $\omega_{\mathrm{c}}$, lies near the resonance frequency, $\omega_{\mathrm{a}}$, of the 87Rb $F=2\rightarrow F^{\prime}=3$ $D_{2}$ transition (wavelength $\lambda=780\,$nm) [35, 36]. The atoms, prepared in the $F=2$ ground state manifold, are exposed to a circularly polarized standing wave pump beam of frequency $\omega_{\mathrm{p}}$ (detuned from the atoms by $\mathrm{\Delta_{pa}}=\omega_{\mathrm{p}}-\omega_{\mathrm{a}}$ and from the cavity by $\mathrm{\Delta_{pc}}=\omega_{\mathrm{p}}-\omega_{\mathrm{c}}$) propagating along the $x$-axis. The circularly polarized pump creates a fictitious magnetic field along $x$ through vector light shifts and defines the atomic quantization axis. The pump drives the $\sigma^{-}$ electric dipole transition of the atoms and its power is linearly ramped up to the desired value in $50\,\mathrm{\mu s}$. The initial ramp optically pumps the atoms into the |$\ket{F=2,m_{F}=-2}$ stretched state, whence the atoms are excited on the cycling transition, thereby restricting spin dynamics to a two-level system. We achieve a maximal vacuum Rabi coupling of $g_{0}=2\pi\times 3.1/\sqrt{2}$ MHz between the $y$-polarized cavity field and the atomic $\sigma^{-}$ transition. The cavity half-linewidth is $\kappa=2\pi\times 0.53\,\mathrm{MHz}$, resulting in a sub-$\mu$s cavity response time defined by $1/\kappa$. Given that the atomic half-linewidth is $\gamma=2\pi\times 3.0\,\mathrm{MHz}$, we achieve a single-atom cooperativity $C=g_{0}^{2}/(2\kappa\gamma)=1.5$ [37]. Because the atom-cavity interaction depends strongly on the spatial structure of the pump and cavity standing waves [38, 39, 40, 41], the tweezers are aligned to be all within the same pump antinode wavefront (so the atoms all experience the same pump Rabi frequency, $\mathrm{\Omega}$) and are positioned at cavity nodes spaced $4.5\lambda$ apart. By spacing the tweezers by a half-integer number of wavelengths, we eliminate biasing in our system due to center-of-mass drift of the tweezer array with respect to the cavity standing wave.
@@
Light inside the cavity creates a potential energy landscape for the atoms through the exchange of cavity photons [27]. For atom $n$ at position $\hat{z}_{n}$, the cavity interaction potential, produced by interference between the pump and cavity fields, is proportional to $S\sin(k\hat{z}_{n})(\hat{c}+\hat{c}^{\dagger})$, where $\hat{c}$ ($\hat{c}^{\dagger}$) is the bosonic annihilation (creation) operator for a cavity photon, $S\simeq\mathrm{\Omega}g_{0}/(2\mathrm{\Delta_{pa}})$ is the coherent scattering rate of a pump photon into the cavity, and $k=2\pi/\lambda$. The strength of the potential thus depends on the atom’s position inside the cavity and on the magnitude of the cavity field. When the pump is sufficiently detuned from the atomic transition and red-detuned from cavity resonance, the cavity field follows the atomic motion such that $\hat{c}+\hat{c}^{\dagger}\propto-S\sum_{m=1}^{N}\sin(k\hat{z}_{m})$. In this adiabatic regime, the effective potential is minimized when $\langle\sin(k\hat{z}_{n})\rangle=\pm 1$ for all $n$. That is, the most favorable configurations of the system are the ones where the atoms self-organize by collectively moving to cavity-field antinodes with the same sign.
@@
The steady state of the optomechanical system is governed by competition between the cavity-mediated interaction and the tweezer potential, which, for tweezers centered at the cavity nodes, is minimized when $\langle\sin(kz_{n})\rangle=0$ for all $n$. This competition gives rise to the emergence of self-organization as a function of $S$, which we tune through $\mathrm{\Omega}$, thus defining a critical pump strength $\mathrm{\Omega_{c}}$. For $\mathrm{\Omega}<\mathrm{\Omega_{c}}$, the tweezers dominate over the cavity field and the atoms remain centered at the cavity nodes (as shown in Fig. 1A(i)). Although the atom array is uncoupled to the cavity when positioned exactly at the nodes, zero-point and thermal fluctuations shift the atom positions, causing the array to emit light incoherently into the cavity. For $\mathrm{\Omega}>\mathrm{\Omega_{c}}$, the negative curvature of the cavity potential overcomes the positive tweezer trap curvature and it becomes energetically favorable for all atoms to move towards antinodes with the same sign and for their emission to interfere constructively. The superradiant field’s phase is determined by the sign of the antinodes chosen by the atoms (Fig. 1A(ii-iii)). For atoms spaced by a half-integer number of wavelengths, adjacent atoms move in opposite directions to achieve self-organization (antiphase motion), in contrast to integer spacing where the atoms move in the same direction (center-of-mass motion). We refer to the collective mode that is linearly coupled to the cavity as the dominant mode.
@@
To detect the complex cavity field $\hat{c}$, we send $y$-polarized emission from one cavity mirror to a heterodyne receiver (see fig. S1) and record the signal for a total time of $250\,\mu\mathrm{s}$, including the $50\,\mu$s pump ramp-up time. We filter the time trace with an adjustable averaging time, ranging from $5\,\mu\mathrm{s}$ to $50\,\mu\mathrm{s}$; we average for a minimum of $5\,\mu$s because increased noise at shorter averaging times compromises the signal-to-noise ratio of the extracted cavity field. The values obtained at each step of the filter are treated as individual data points for $\hat{c}$. We determine the angle $\theta$ by referencing the phase of the distribution of $\hat{c}$ to that obtained with all atoms placed at antinodes of the same sign. We then project each value of $\hat{c}$ onto the axis defined by $\theta$ to obtain $c_{\mathrm{proj}}$ (see materials and methods and fig. S2). Our measured $c_{\mathrm{proj}}$ is scaled such that $c_{\mathrm{proj}}^{2}$ is the average photon number in the cavity.
@@
However, thermal fluctuations as well as motion of the atoms perpendicular to the cavity axis are relevant for a quantitative description of the experiment. By means of a more refined optomechanical model [45], the critical pump strength is found to be
@@
where $T$ is the atomic motional temperature and the temperature-dependent factor $\varepsilon(T)$ is defined in the supplementary text. The dependencies of $\mathrm{\Omega_{c}}$ on $N$ and $\mathrm{\Delta_{pa}}$ are studied in this work (see Fig. 2 and fig. S4). Equation (4) includes two different thermal effects on the critical pump strength. First, taking an average over the thermal motional shift of the atoms yields a net shift of the cavity resonance $\Delta_{\text{pc}}(T)\equiv\langle\tilde{\Delta}_{\text{pc}}\rangle$. Second, thermal fluctuations allow the atoms to probe the non-linearity of the effective mechanical potential in $\hat{\Theta}^{2}$. This leads to an increase in the critical pump strength proportional to the factor $1/\varepsilon(T)$ in Equation (4). Using a temperature of $35\pm 10\,\mu$K, determined from an independent measurement [38], Equation (4) predicts $\mathrm{\Omega_{c}}=2\pi\times(26.5\pm 2.5)\,$MHz for the parameter settings in Fig. 1B.
@@
where $B$ and $D$ are parameter-dependent constants and $z_{\mathrm{dom}}\equiv\langle\hat{z}_{\mathrm{dom}}\rangle$. According to Landau’s theory, a second order phase transition occurs at $B=0$. Because $c_{\mathrm{proj}}$ is proportional to $z_{\mathrm{dom}}$ to leading order, we fit the observed distributions of $c_{\mathrm{proj}}$ to Equation (5) and extract the coefficient $B$ to determine where $\mathrm{\Omega_{c}}$ occurs. Analyzing the distributions of $c_{\mathrm{proj}}$, we observe that the fits to the Boltzmann distribution match the data well (see fig. S5 and supplementary text) even though the system is driven and dissipative. From the data in Fig. 1B, we extract an experimental $\mathrm{\Omega_{c}}=2\pi\times(33.1^{+2.0_{\mathrm{syst}}}_{-2.8_{\mathrm{syst}}}\pm 1.6_{\mathrm{stat}})\,\mathrm{MHz}$. This measurement is slightly above the prediction of Equation (4) for our independent measurement of the initial temperature of the trapped atoms. The slight discrepancy may be caused by heating of the trapped atoms during the experiment, owing to light scattering. Nonetheless, when statistical and systematic errors (primarily resulting from uncertainty in our calibration of $\Omega$ [45]) are taken into consideration, the measurement falls within the error of the prediction of Equation (4) for our independent temperature measurement.
Figure 2: Dependence of the critical point on atom number. (A) Bifurcation data for $N=10$ (circles), $N=16$ (squares), and $N=22$ (triangles). Shaded areas show the width of the distribution at half maximum while markers show the maxima of the fitted Boltzmann distrbution. (B) Experimentally measured $\mathrm{\Omega_{c}}$ (circles) for different $N$ plotted on a logarithmic scale and the theoretical prediction of Equation (4), based on an independent temperature measurement of $T=35\pm 10\,\mu$K (black dashed line and light blue shaded area). Error bars represent statistical errors from fits of the parameter $B$ and drift of the pump power over the course of the experiment, estimated to be 10%. The reported values of $\Omega_{\mathrm{c}}$ account for systematic errors due to Gaussian detector noise and the Gaussian intensity profile of the pump beam. We note that there is an additional systematic error of +6.0% and -8.5% on our calibration of $\Omega$ (see supplementary text for details). All data in figure are taken at $\mathrm{\Delta_{pa}}=-2\pi\times 80\,\mathrm{MHz}$ and $\mathrm{\Delta_{pc}}=-2\pi\times 1.9\,\mathrm{MHz}$.

### v2

@@ Abstract
Increasing the number of particles in a system often leads to qualitative changes in its properties, such as breaking of symmetries and the appearance of phase transitions. This renders a macroscopic system fundamentally different from its individual microscopic constituents. Lying between these extremes, mesoscopic systems exhibit microscopic fluctuations that influence behavior on longer length scales, leading to critical phenomena and dynamics. Therefore, tracing the properties of well-controlled mesoscopic systems can help bridge the gap between an exact description of few-body microscopic systems and the emergent description of many-body systems. Here, we explore mesoscopic signatures of an optomechanical self-organization phase transition using arrays of cold atoms inside an optical cavity. By precisely engineering atom-cavity interactions, we reveal how critical behavior depends on atom number, identify characteristic dynamical behaviors in the self-organized regime, and observe a finite optomechanical susceptibility at the critical point. These findings advance our understanding of particle-number- and time-resolved properties of phase transitions in mesoscopic systems.
@@
The scale of a physical system plays a key role in determining its material properties. With increasing size comes increasing complexity, and complexity leads to the emergence of macroscopic phenomena that, often, cannot be extrapolated from microscopic behaviors [1]. A well-known example of this paradigm is the appearance of phase transitions in the thermodynamic limit, i.e. as a system’s particle number tends to infinity. In between the microscopic and macroscopic extremes lies the mesoscopic regime. Mesoscopic systems not only exhibit characteristics of macroscopic systems, but also host a variety of rich properties themselves. In systems of intermediate size, assumptions of local equilibrium break down and internal microstructures may affect properties on longer length scales [2]. The fact that microscopic fluctuations have a measurable influence on mesoscopic scales also makes mesoscopic systems an interesting playground for the study of critical phenomena and phase transitions [3, 4, 5].
@@
Technical advances in engineering, probing and measuring mesoscopic systems have propelled experiments in a variety of physical platforms. Notable examples include solid state systems [6, 7], nanostructured [8] and nanomechanical systems [9, 10, 11, 12], nanomagnets [13, 14], and cold atoms and ions [15, 16, 17, 18, 19, 20, 21]. Breakthroughs in the control of atoms trapped in optical tweezer arrays [22, 23] represent a leap forward in precisely constructing mesoscopic, ultracold atomic systems. We show that deterministic control over particle number and interaction strength can be achieved with an optical tweezer array inside a standing-wave optical cavity. This system uniquely lends itself to the detailed study of optomechanics in mesoscopic samples.
@@ I Cavity-Mediated interactions and Self-organization
Figure 1: Cavity-induced self-organization. a, Schematic of the experimental setup. Optically tweezed atoms are placed at the nodes of the cavity field at half-integer wavelength spacing (1.5 $\lambda$ shown in figure for convenience). The atoms are driven by a pair of circularly polarized counter-propagating pump beams. The transmitted cavity field is detected with a heterodyne receiver. Insets (i-iii) illustrate self-organization. (i) At pump strengths below the critical point, the atoms remain centered at the cavity nodes and on average emit zero field. Above the critical point, atoms self-organize by moving toward either the (ii) positive (red) or (iii) negative (blue) antinodes. b, Probability distributions of cavity field projections from repeated measurements. The results shown are based on data collected with $N=20$, $\mathrm{\Delta_{pc}}=-2\pi\times 1.9\,\mathrm{MHz}$ and $\mathrm{\Delta_{pa}}=-2\pi\times 80\,\mathrm{MHz}$. The field distributions are plotted for different values of $\mathrm{\Omega}$. The data are averaged over $5\,\mu\mathrm{s}$ intervals and are sampled from the first $100\,\mu\mathrm{s}$ after the pump ramp-up has finished. The labels (i), (ii), and (iii) correspond to the drawings in (a). The white dashed line indicates the critical point fitted from the data.
@@
To realize mesoscopic self-organization, we use the experimental setup schematically shown in Fig. 1a. It contains a one-dimensional tweezer array into which we load a deterministic number $N$ of 87Rb atoms. The array is aligned along the axis of a high-finesse Fabry-Pérot optical cavity (the $z$-axis). The cavity is placed in-vacuum and has a TEM00 mode whose frequency, $\omega_{\mathrm{c}}$, lies near the resonance frequency, $\omega_{\mathrm{a}}$, of the 87Rb $F=2\rightarrow F^{\prime}=3$ $D_{2}$ transition (wavelength $\lambda=780\,$nm) [33, 34]. The atoms, prepared in the $F=2$ ground state manifold, are exposed to a circularly polarized standing wave pump beam of frequency $\omega_{\mathrm{p}}$ (detuned from the atoms by $\mathrm{\Delta_{pa}}=\omega_{\mathrm{p}}-\omega_{\mathrm{a}}$ and from the cavity by $\mathrm{\Delta_{pc}}=\omega_{\mathrm{p}}-\omega_{\mathrm{c}}$) propagating along the $x$-axis. The circularly polarized pump creates a fictitious magnetic field along $x$ through vector light shifts and defines the atomic quantization axis. The pump drives the $\sigma^{-}$ electric dipole transition of the atoms and its power is linearly ramped up to the desired value in $50\,\mathrm{\mu s}$. The initial ramp optically pumps the atoms into the $\lvert F=2,m_{F}=-2\rangle$ stretched state, whence the atoms are excited on the cycling transition, thereby ensuring that no dynamics among the $\lvert F=2,m_{F}\rangle$ states interfere with the location of the critical point for self-organization. We achieve a maximal vacuum Rabi coupling of $g_{0}=2\pi\times 3.1/\sqrt{2}$ MHz between the $y$-polarized cavity field and the atomic $\sigma^{-}$ transition. The cavity half-linewidth is $\kappa=2\pi\times 0.53\,\mathrm{MHz}$, resulting in a sub-$\mu$s cavity response time defined by $1/\kappa$. Given that the atomic half-linewidth is $\gamma=2\pi\times 3.0\,\mathrm{MHz}$, we achieve a single-atom cooperativity $C=g_{0}^{2}/(2\kappa\gamma)=1.5$ [35]. Because the atom-cavity interaction depends strongly on the spatial structure of the pump and cavity standing waves [36, 37, 38, 39], the tweezers are aligned to be all within the same pump antinode wavefront (so the atoms all experience the same pump Rabi frequency, $\mathrm{\Omega}$) and are positioned at cavity nodes spaced $4.5\,\lambda$ apart. By spacing the tweezers by a half-integer number of wavelengths, we eliminate biasing in our system due to center-of-mass drift of the tweezer array with respect to the cavity standing wave.
@@
Light inside the cavity creates a potential energy landscape for the atoms through the exchange of cavity photons [25]. For atom $n$ at position $\hat{z}_{n}$, the cavity interaction potential, produced by interference between the pump and cavity fields, is proportional to $S\sin(k\hat{z}_{n})(\hat{c}+\hat{c}^{\dagger})$, where $\hat{c}$ ($\hat{c}^{\dagger}$) is the bosonic annihilation (creation) operator for a cavity photon, $S\simeq\mathrm{\Omega}g_{0}/(2\mathrm{\Delta_{pa}})$ is the coherent scattering rate of a pump photon into the cavity, and $k=2\pi/\lambda$. The strength of the potential thus depends on the atom’s position inside the cavity and on the magnitude of the cavity field. When the pump is sufficiently detuned from the atomic transition and red-detuned from cavity resonance, the cavity field follows the atomic motion such that $\hat{c}+\hat{c}^{\dagger}\propto-S\sum_{n=1}^{N}\sin(k\hat{z}_{n})$. In this adiabatic regime, the effective potential is minimized when $\langle\sin(k\hat{z}_{n})\rangle=\pm 1$ for all $n$. That is, the most favorable configurations of the system are the ones where the atoms self-organize by collectively moving to cavity-field antinodes with the same sign.
@@
The steady state of the optomechanical system is governed by competition between the cavity-mediated interaction and the tweezer potential, which, for tweezers centered at the cavity nodes, is minimized when $\langle\sin(k\hat{z}_{n})\rangle=0$ for all $n$. This competition gives rise to the emergence of self-organization as a function of $S$, which we tune through $\mathrm{\Omega}$, thus defining a critical pump strength $\mathrm{\Omega_{c}}$. For $\mathrm{\Omega}<\mathrm{\Omega_{c}}$, the tweezers dominate over the cavity field and the atoms remain centered at the cavity nodes (as shown in Fig. 1a(i)). Although the atom array is uncoupled from the cavity when positioned exactly at the nodes, zero-point and thermal fluctuations shift the atom positions, causing the array to emit light incoherently into the cavity. For $\mathrm{\Omega}>\mathrm{\Omega_{c}}$, the negative curvature of the cavity potential overcomes the positive tweezer trap curvature and it becomes energetically favorable for all atoms to move towards antinodes with the same sign and for their emission to interfere constructively. The superradiant field’s phase is determined by the sign of the antinodes chosen by the atoms (Fig. 1a(ii-iii)). For atoms spaced by a half-integer number of wavelengths, adjacent atoms move in opposite directions to achieve self-organization (antiphase motion). We refer to the collective mode that is linearly coupled to the cavity as the dominant mode.
@@
To detect the complex cavity field $\hat{c}$, we send $y$-polarized emission from one cavity mirror to a heterodyne receiver (see Extended Data Fig. 1) and record the signal for a total time of $250\,\mu\mathrm{s}$, including the $50\,\mu$s pump ramp-up time. We filter the time trace with an adjustable averaging time, ranging from $5\,\mu\mathrm{s}$ to $50\,\mu\mathrm{s}$; we average for a minimum of $5\,\mu$s because increased noise at shorter averaging times compromises the signal-to-noise ratio of the extracted cavity field. The values obtained at each step of the filter are treated as individual data points for $\hat{c}$. We determine the angle $\theta$ by referencing the phase of the distribution of $\hat{c}$ to that obtained with all atoms placed at antinodes of the same sign. We then project each value of $\hat{c}$ onto the axis defined by $\theta$ to obtain $c_{\mathrm{proj}}$ (see Methods and Extended Data Fig. 2). Our measured $c_{\mathrm{proj}}$ is scaled such that $c_{\mathrm{proj}}^{2}$ is the average photon number in the cavity.
@@
However, thermal fluctuations as well as motion of the atoms perpendicular to the cavity axis are relevant for a quantitative description of the experiment. By means of a more refined optomechanical model (see Supplementary Information Section II), the critical pump strength is found to be
@@
where $T$ is the atomic motional temperature and the temperature-dependent factor $\varepsilon(T)$ is defined in Supplementary Equation (S29). The dependencies of $\mathrm{\Omega_{c}}$ on $N$ and $\mathrm{\Delta_{pa}}$ are studied in this work (see Fig. 2 and Extended Data Fig. 3). Equation (4) includes two different thermal effects on the critical pump strength. First, taking an average over the thermal motional shift of the atoms yields a net shift of the cavity resonance $\Delta_{\text{pc}}(T)\equiv\langle\tilde{\Delta}_{\text{pc}}\rangle$. Second, thermal fluctuations allow the atoms to probe the non-linearity of the effective mechanical potential in $\hat{\Theta}^{2}$. This leads to an increase in the critical pump strength proportional to the factor $1/\varepsilon(T)$ in Equation (4). Using a temperature of $35\pm 10\,\mu$K, determined from an independent measurement [36], Equation (4) predicts $\mathrm{\Omega_{c}}=2\pi\times(26.5\pm 2.5)\,$MHz for the parameter settings in Fig. 1b.
@@
where $B$ and $D$ are parameter-dependent constants and $z_{\mathrm{dom}}\equiv\langle\hat{z}_{\mathrm{dom}}\rangle$. According to Landau’s theory, a second order phase transition occurs at $B=0$. Because $c_{\mathrm{proj}}$ is proportional to $z_{\mathrm{dom}}$ to leading order, we fit the observed distributions of $c_{\mathrm{proj}}$ to Equation (5) and extract the coefficient $B$ to determine where $\mathrm{\Omega_{c}}$ occurs. Analyzing the distributions of $c_{\mathrm{proj}}$, we observe that the fits to the Boltzmann distribution match the data well (see Extended Data Fig. 4 and Methods) even though the system is driven and dissipative. From the data in Fig. 1b, we extract an experimental $\mathrm{\Omega_{c}}=2\pi\times(33.1^{+2.0_{\mathrm{syst}}}_{-2.8_{\mathrm{syst}}}\pm 1.6_{\mathrm{stat}})\,\mathrm{MHz}$. This measurement is slightly above the prediction of Equation (4) for our independent measurement of the initial temperature of the trapped atoms, though within error. The slight discrepancy may be caused by heating of the trapped atoms during the experiment, owing to light scattering. In future work, it may be possible to monitor the atoms’ temperature during the experiment by observing sideband asymmetry in the heterodyne power spectrum.
Figure 2: Dependence of the critical point on atom number. a, Bifurcation data for $N=10$ (circles), $N=16$ (squares), and $N=22$ (triangles). Shaded areas show the width of the distribution at half maximum while markers show the maxima of the fitted Boltzmann distrbution. b, Experimentally measured $\mathrm{\Omega_{c}}$ (circles) for different $N$ plotted on a logarithmic scale and the theoretical prediction of Equation (4), based on an independent temperature measurement of $T=35\pm 10\,\mu$K (black dashed line and light blue shaded area). Error bars represent standard deviations resulting from statistical errors on fits of the parameter $B$ and drift of the pump power over the course of the experiment, estimated to be 10%. The reported values of $\Omega_{\mathrm{c}}$ account for systematic errors due to Gaussian detector noise and the Gaussian intensity profile of the pump beam. We note that there is an additional systematic error of +6.0% and -8.5% on our calibration of $\Omega$ (see Methods for details). All data in figure are taken at $\mathrm{\Delta_{pa}}=-2\pi\times 80\,\mathrm{MHz}$ and $\mathrm{\Delta_{pc}}=-2\pi\times 1.9\,\mathrm{MHz}$. The value of $\mathrm{\Delta_{pa}}$ was chosen to avoid strong atom number-dependence of the atom-induced dispersive shift on the cavity resonance frequency. We also chose the bare cavity resonance frequency to be sufficiently larger than the pump frequency to avoid instabilities arising from the dispersive shift bringing the pump blue-detuned of the cavity.